# thanks: corresponding author

### v1

@@ 2.2 Formal Equivalence Proof Through Operator Inversion, Mercer’s Theorem and Regularization
2.2.1 Operator Perspective in SVMs and Quantum Mechanics
@@
SVMKernel Matrix $\mathbf{Q}$Operator InversionSolution $\boldsymbol{\alpha}$Quantum MechanicsHamiltonian $\hat{H}$perator InversionGreen’s Function $G$ Operator Inversion Equivalence
Figure 1: Diagram illustrating the equivalence through operator inversion between SVMs and quantum mechanics. Both require inverting a (positive) operator to obtain the final solution (either $\boldsymbol{\alpha}$ in SVM or $G$ in quantum mechanics).
@@
In quantum mechanics or PDE settings, the Green’s function $G$ similarly emerges from inverting a self-adjoint operator, such as $\hat{H}-E$ for the Hamiltonian. Figure 1 illustrates this structural equivalence: SVMs invert $\mathbf{Q}$ to obtain $\boldsymbol{\alpha}$, while quantum/PDE systems invert their Hamiltonian or differential operator to obtain the Green’s function $G$.
As mentioned previously, a rigorous equivalence between kernels and Green’s functions demands the following conditions are met:
@@
2.2.2 Eigenfunction Expansion and Mercer’s Theorem vs. PDE Expansion
Mercer’s theorem [13] originally emerged from the study of integral equations, providing a rigorous mathematical foundation for symmetric positive-definite kernels. The theorem’s relevance for machine learning was established through its application to kernel methods [11, 3]. This connection enabled the use of kernels in SVMs and beyond, thereby pushing forward the field of machine learning by allowing it to embed data into high-dimensional feature spaces without explicitly constructing these mappings (the so-called kernel trick).
As for Green’s functions, the expansion of partial differential equations (PDEs) in terms of eigenfunctions [14] is a cornerstone in classical analysis. Mercer’s theorem and PDE expansion both involve the decomposition of operators into eigenfunctions, with Mercer kernels providing a spectral decomposition similar to the eigenfunction solutions of PDEs.
Both the Mercer expansion for a positive semi-definite kernel and the eigenfunction expansion of a self-adjoint operator’s Green’s function decompose the respective objects into a sum of rank-one contributions. This similarity underscores the fact that in both cases the underlying eigenvalues and eigenfunctions govern the behavior of the system. Here, we use this observation to build a unifying framework that connects SVM kernel methods with operator inversion in quantum mechanics.
Spectral or Eigenfunction Decompositions
SVM Kernel Operator. According to Mercer’s theorem and standard linear algebra, a symmetric positive semi-definite kernel $K$ can always be expressed as a spectral decomposition:
@@
Green’s Function Operator. In quantum mechanics or PDEs, the resolvent operator $(\hat{H}-E)^{-1}$ associated with a Hamiltonian $\hat{H}$ can similarly be expressed as:
@@
Structural Similarity. Both $K$ and $G$ share a common structure: they are sums of rank-1 outer products weighted by spectral coefficients. The kernel $K$ corresponds to a weighted inner product in a high-dimensional feature space, while $G$ represents the propagator in a quantum or PDE system. Under suitable conditions, namely when the domain, boundary conditions, and positivity constraints of the problem coincide, it is possible to establish an isomorphism between the eigenfunction decompositions of the SVM kernel and the Green’s function. In this sense, one may regard $K$ and $G$ as equivalent representations of the same underlying operator inversion (see, e.g., [13, 3, 7] for related discussions).
@@
In machine learning, PSD kernel functions serve as the cornerstone of SVMs and kernel methods. By implicitly mapping input data into high-dimensional feature spaces, kernel functions enable nonlinear separations while satisfying Mercer’s theorem to ensure positive definiteness [3, 11].
Kernel functions can approximate the Green’s functions of PDEs, providing numerical tools for physical modeling. For example: Eq. (11) is analogous to the spectral decomposition of Green’s functions in PDEs, as in Eq. (12). While the mathematical structures are similar, their application backgrounds differ. Kernel functions are primarily used to capture data relationships in high-dimensional spaces, while Green’s functions are fundamental solutions to physical equations.
Positivity and Boundary Conditions.
Mercer’s Theorem for SVM. For $K(\mathbf{x},\mathbf{x}^{\prime})$ to be a valid SVM kernel on domain $\Omega$, it must be continuous, symmetric, and positive semi-definite (PSD). This ensures a convex SVM dual.
Green’s Function PDE Setup. For $G(\mathbf{x},\mathbf{x}^{\prime})$ to serve as an inverse of $(\hat{H}-E)$, we typically require self-adjointness and positivity. Boundary conditions in PDE contexts (Dirichlet, Neumann, etc.) are responsible for the exact form of $G$. When the domain, boundary conditions, and positivity constraints of an SVM kernel and the corresponding Green’s function are aligned, one can formally establish an equivalence between the two (see, e.g., [13, 7]). Although this is subject to specific assumptions, it serves as a useful guiding principle for designing kernels that reflect the underlying physics.
@@
does not universally guarantee positive semi-definiteness. Although included as a standard option in many SVM implementations, the sigmoid kernel only meets Mercer’s condition under restrictive parameter regimes (e.g., specific ranges of $\gamma>0$ and $r<0$). Outside these regimes, the kernel matrix may lose its PSD property, leading to non-convex optimization or instability in training.
This sensitivity to parameter choices underlines why the sigmoid kernel is less used in practice than kernels like the RBF that satisfy Mercer’s theorem. This also complicates its use in drawing rigorous parallels between SVM kernels and Green’s functions, since the latter rely on strict positivity and well-defined spectral properties. As a result, while attempting to establish a precise equivalence between SVM kernels and Green’s functions, one must either operate within the safe parameter bounds for the sigmoid kernel or consider alternative kernel choices that inherently satisfy the necessary positivity conditions.
2.2.3 Margin Maximization and Action Extremization
Both SVMs and quantum mechanics involve optimizing a functional. For SVMs, one optimizes the margin through an objective that depends on a kernel $K$. In quantum and Langrangian mechanics, one extremizes the action $S$, which often leads to Green’s functions through operator inversion.
In SVM formulations, the decision hyperplane, which separates classes or fits a regression function, arises from the corresponding optimization problem, which is often solved via its dual formulation:
@@ 2.2.3 Margin Maximization and Action Extremization
We can draw a parallel between SVM margin maximization and action extremization in physics. In SVMs, one seeks to maximize the margin by solving a variational problem in a high-dimensional feature space, which leads to a solution expressible as a weighted sum of kernel evaluations. In a similar manner, physical systems are often characterized by the principle of least action, where the trajectory or state of the system is obtained by extremizing a Lagrangian functional. In both cases, the solution is determined by finding a stationary point of an objective functional, and - in view of the representer theorem - the optimal solution can be represented as a finite linear combination of basis functions (kernel functions in SVMs, or Green’s functions in physical systems).
@@
For this reason, we may interpret a physical system’s trajectory or field configuration (obtained via the Lagrangian) as analogous to the SVM’s decision hyperplane. Determination of the optimal trajectory through a variational principle resembles selecting the optimal hyperplane in SVMs via margin maximization. This perspective allows us to recast certain equations of motion (e.g., in regression or classification tasks) within an SVM framework, particularly when the operator or Green’s function governing the physical system is used as the SVM kernel. In the following sections, we expand on these ideas, illustrating how diverse physical quantities can be modeled through a kernel-based approach once the underlying action principle and the associated Green’s function are identified.
In addition to the structural equivalence established by operator inversion, the representer theorem guarantees that the solution to a regularized risk minimization problem can be expressed as a finite sum of kernel (or Green’s function) evaluations. This further solidifies the connection between SVM methods and physical systems governed by Green’s functions.
2.2.4 Link to Equivalence of Regularization Solutions and Green’s Functions
While the operator inversion perspective establishes a structural similarity between SVM kernels and Green’s functions, it primarily highlights that both arise from inverting a positive (semi-definite) operator. However, this conceptual connection alone does not fully detail how solutions manifest in practice. We discuss the equivalence of regularization solutions to address this gap by providing a constructive framework that complements and extends the operator inversion argument.
A further result, given in Theorem 2.1, states that regularization solutions in PDE/inverse problems can be expanded as sums of Green’s functions at data points, mirroring how SVM solutions expand in sums of kernel evaluations. Therefore, the following two points are complementary:
  • Recognition of the fact that SVM dual solving $\mathbf{Q}\,\boldsymbol{\alpha}=\cdots$ is analogous to PDE-based Green’s function solving $(\hat{H}-E)\,G=\delta$.
  • Equivalence of regularization solutions (Theorem 2.1), which ensures that PDE-based minimizers must lie in the span of Green’s functions. This parallels the representer theorem in kernel-based ML and clarifies how both are expansions in the same operator’s inverse (assuming positivity, boundary conditions, etc.).
To summarize, while the operator inversion perspective offers a conceptual basis for the equivalence between SVM kernels and Green’s functions, Theorem 2.1 and its discussion supplies the detailed constructive link. They validate that solutions obtained via regularization in PDE contexts can be expressed explicitly as expansions in Green’s functions, analogous to SVM solutions as sums of kernel evaluations. This constructive equivalence reinforces the operator inversion argument by incorporating spectral positivity, boundary constraints, and functional expansions.
2.3 Formulating Physical Problems as Optimization Problems
Building on the established connection between operator inversion and the equivalence of regularization solutions with Green’s functions, we now demonstrate how many physical problems can be reformulated as optimization tasks.
Many problems in physics that involve Green’s functions, such as solving PDEs or describing quantum systems, can be reformulated as optimization or variational problems. Consider, for instance, a physical system described by a linear differential operator $P$ (e.g., the Laplace operator). The goal is to determine the physical field $u(x)$ that satisfies certain constraints or fits observed data.
By the representer theorem [3], solutions to regularization problems in a reproducing kernel Hilbert space can be expressed as finite linear combinations of kernel (or Green’s function) evaluations at the data points. With this, many physical problems can be cast as minimization tasks over function spaces.
Our particular example can be cast as the following minimization problem:
$f^{*}=\arg\min_{f\in H}\left(\sum_{i=1}^{N}V(y_{i},f(x_{i}))+\lambda\|Pf\|_{L^{2}}^{2}\right)$ (15)
where:
  • $V(y_{i},f(x_{i}))$ is the loss function measuring the discrepancy between the observed value $y_{i}$ and the predicted value $f(x_{i})$,
  • $\|Pf\|_{L^{2}}^{2}$ is a regularization term controlling the complexity of $f$, and
  • $\lambda$ is a regularization parameter balancing the trade-off between fitting the data and controlling complexity.
The loss term $V(y_{i},f(x_{i}))$ in the functional resembles the role of a Lagrangian density in physics, where discrepancies between $f(x_{i})$ and $y_{i}$ contribute to the overall "action" that the system seeks to extremize. In this way, the optimization functional parallels a variational principle: just as the Lagrangian integrates over fields and their dynamics to yield equations of motion, the loss term adds up discrepancies to guide the selection of an optimal function $f^{*}$. This perspective demonstrates how minimizing such a functional in the context of SVMs or regularization problems is akin to finding a stationary action in physics, thereby providing a direct link between the loss function and physical principles.
On the other hand, this optimization formulation is analogous to finding the optimal decision hyperplane in SVMs, where the goal is to minimize a regularized risk functional. Just as the SVM decision function can be expressed as a sum of kernel evaluations weighted by optimal coefficients (derived from solving a dual problem), the solution $f^{*}$ here is represented as a combination of Green’s functions centered at the data points. This parallel suggests that techniques developed in SVMs can be directly applied to solve complex physical problems.
The Green’s function $G(x,x^{\prime})$, defined as the inverse of the operator $P$, plays a fundamental role here. By minimizing the above functional, the solution $f^{*}(x)$ can often be expressed as a linear combination of Green’s functions centered at the data points.
In quantum mechanics, the wavefunction $\psi(x)$ is typically obtained by solving the Schrödinger equation, which is inherently a variational problem. Take for example
$\psi^{*}=\arg\min_{\psi}\langle\psi|\hat{H}|\psi\rangle$ (16)
subject to $\|\psi\|=1$, where $\hat{H}$ is the Hamiltonian operator. This optimization minimizes the energy functional and is structurally similar to SVM optimization problems, with the Hilbert space playing the role of the feature space.
In both cases, the minimizer can be expressed in terms of Green’s functions, highlighting the intrinsic connection between the two fields.
As discussed above, the link between SVM kernels and Green’s functions becomes clearer through the perspective of operator inversion. Both kernel functions $K(x,x^{\prime})$ in machine learning and Green’s functions $G(x,x^{\prime})$ in physics are solutions to operator-based equations. For SVMs, $K(x,x^{\prime})$ corresponds to the solution of linear systems involving the kernel matrix $\mathbf{K}$, which is derived from the inner product in the feature space. For physics, $G(x,x^{\prime})$ is the inverse of a differential operator, describing how a system responds to a localized perturbation.
In both contexts, the solution can be represented as an inner product in a high-dimensional space, making the representer theorem directly applicable. This mathematical equivalence enables us to replace $K$ with $G$ when solving regularization problems.
We now give a proof in Theorem 2.1 that $f^{*}$ as defined in Eq. (20) can be expressed as a superposition of Green’s functions centered as the datapoints.
Theorem 2.1.
The function $f^{*}$ minimizing the regularization functional
$\mathcal{R}[f]=\sum_{i=1}^{N}V(y_{i},f(x_{i}))+\lambda\|Pf\|_{L^{2}}^{2},$ (17)
where $P$ is a linear differential operator, can be expressed as a superposition of Green’s functions centered at the data points:
$f^{*}(x)=\sum_{i=1}^{N}\alpha_{i}G(x,x_{i}),$ (18)
highlighting the equivalence between the regularization solution and the Green’s function.
Proof.
By the representer theorem [3] and the properties of the reproducing kernel Hilbert space (RKHS) associated with the operator $P$, the minimizer of the functional $\mathcal{R}[f]$ lies in the span of the Green’s functions $G(x,x_{i})$. Specifically, the Green’s function $G(x,x^{\prime})$ of $P^{*}P$, where $P^{*}$ is the adjoint operator of $P$, satisfies:
$P^{*}PG(x,x^{\prime})=\delta(x-x^{\prime}),$ (19)
with $\delta(x-x^{\prime})$ being the Dirac delta function.
Substituting this representation into the functional, it follows that the minimizer $f^{*}(x)$ can be written as:
$f^{*}(x)=\sum_{i=1}^{N}\alpha_{i}G(x,x_{i}),$ (20)
where $\alpha_{i}$ are coefficients determined by the optimization process. ∎
In computational physics, the solution in Eq. (20) arises as the minimizer of a regularization problem involving a linear differential operator $P$. Specifically, we consider a physical system governed by $P$ and suppose we seek a function $f(x)$ that fits observed data $\{(x_{i},y_{i})\}_{i=1}^{N}$ while controlling its complexity. This leads to the minimization problem:
$\min_{f\in H}\left(\sum_{i=1}^{N}V(y_{i},f(x_{i}))+\lambda\|Pf\|_{L^{2}}^{2}\right).$
By the representer theorem, the solution lies in the span of the Green’s functions $G(x,x_{i})$ associated with the operator $P^{*}P$. For example, in the case of the Poisson equation
$-\Delta u(x)=\rho(x),\quad x\in\Omega,\quad u|_{\partial\Omega}=0,$
the Green’s function $G(x,x^{\prime})$ satisfies
$-\Delta_{x}G(x,x^{\prime})=\delta(x-x^{\prime}),\quad G(x,x^{\prime})|_{x\in\partial\Omega}=0.$
Then, the regularized solution for approximating $u(x)$ from data $u(x_{i})\approx y_{i}$ can be written as
$u^{*}(x)=\sum_{i=1}^{N}\alpha_{i}G(x,x_{i}),$
which explicitly uses the Green’s function of the Laplacian to interpolate and regularize the solution.
In the context of SVMs, the analogous expression
$f^{*}(x)=\sum_{i=1}^{N}\alpha_{i}K(x,x_{i})$ (21)
is derived from solving a regularized optimization problem in a RKHS. Here, the kernel function $K(x,x_{i})$ serves a role similar to that of a Green’s function, but in the machine learning setting, it implicitly defines a mapping into a high-dimensional feature space.
More concretely, consider a linear differential operator $P$ that maps input data $x$ to a feature space through a transformation $\phi(x)$, i.e.,
$\phi(x)=P\,x,$
where $P$ acts as the linear mapping (or feature extraction operator). The kernel function is then defined as the inner product in the feature space:
$K(x,x^{\prime})=\langle\phi(x),\phi(x^{\prime})\rangle.$
Using the representer theorem in this RKHS setting, the solution to a regularized risk minimization problem is given by Eq. (21), which mirrors the structure of the physics solution with $G(x,x_{i})$. Here, the operator $P$ used to map the input into the feature space is analogous to the differential operator in the physics context. The coefficients $\alpha_{i}$ are obtained by solving a dual optimization problem, similar to how they are determined in the PDE regularization setting.
Thus, the expressions in Eq. (20) in physics and Eq. (21) in SVMs both represent solutions expanded in terms of Green’s functions and kernel functions, respectively, highlighting their equivalence under suitable conditions.
2.3.1 Algorithm: Solving Regularization Problems Using Kernels
Building on Theorem 2.1 and the operator inversion perspective, we present a general algorithm for solving regularization problems by constructing a kernel from a known self-adjoint differential operator $P$. This approach is widely applicable in computational physics, where $P$ is derived from the governing equations of the system.
Algorithm 1 Solving Regularization Problems with Operator $P$
0:  Data $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, differential operator $P$, regularization parameter $\lambda$
0:  Function $f^{*}(\mathbf{x})$
1:  Compute the Green’s function $G(\mathbf{x},\mathbf{x}^{\prime})$ satisfying $P^{*}PG=\delta(\mathbf{x}-\mathbf{x}^{\prime})$.
2:  Form the kernel matrix $\mathbf{K}$ with elements $K_{ij}=G(\mathbf{x}_{i},\mathbf{x}_{j})$.
3:  Solve the linear system:
$(\mathbf{K}+\lambda\mathbf{I})\boldsymbol{\alpha}=\mathbf{y},$
where $\boldsymbol{\alpha}=[\alpha_{1},\dots,\alpha_{N}]^{\top}$, $\mathbf{I}$ is the identity matrix, and $\mathbf{y}=[y_{1},\dots,y_{N}]^{\top}$.
4:  Construct $f^{*}(\mathbf{x})=\sum_{i=1}^{N}\alpha_{i}G(\mathbf{x},\mathbf{x}_{i})$.
This algorithm outlines a general procedure. It begins by computing the Green’s function $G$ for the given self-adjoint operator $P$, establishing the connection to the underlying physical model. Then, using $G$, a kernel matrix $\mathbf{K}$ is formed, and a regularized linear system is solved to obtain coefficients $\boldsymbol{\alpha}$. And finally, the solution $f^{*}(\mathbf{x})$ is expressed as a linear combination of Green’s functions centered at the data points.
In many practical applications, especially within the scope of Section 3, the Green’s function $G(\mathbf{x},\mathbf{x}^{\prime})$ corresponding to the physical process is already known from theory or prior computation. In such cases, the initial step of computing $G$ from $P$ can be skipped. The algorithm then simplifies to directly using the known $G$ to form the kernel matrix, solve for $\boldsymbol{\alpha}$, and construct the regression function. This special case is a direct application of the general procedure when the self-adjoint operator $P$ is implicit in the known Green’s function.
3 Applications: SVM Regression in Physical Systems
In this section, we consider several aspects of physical systems and how particular kernels perform for these systems. All the examples involve regression using SVMs. SVMs’ use for regression is standard [30], with established kernels such as RBF, linear, sigmoid or polynomial [3]. One of the example aspects considered is conductivity, for which we will be using the Kubo formula. Physical modeling of conductivity via the Kubo formula is well-documented; see, for instance, Mahan [31]. In electron-proton scattering, we apply the Born approximation, the description of which can be readily found in standard quantum mechanics references [33].

### v2

@@ 3.2 Other forms of equivalence
3.2.1 Operator inversion
@@
Support Vector Machine (SVM)Kernel Matrix $\mathbf{Q}$Operator InversionSolution $\boldsymbol{\alpha}$Quantum MechanicsHamiltonian $\hat{H}$perator InversionGreen’s Function $G$ Operator Inversion Equivalence
Figure 1: Schematic illustration of the operator inversion equivalence between SVM and quantum mechanics.
@@
In quantum mechanics or PDE settings, Green’s function $G$ similarly emerges from inverting a self-adjoint operator, such as $\hat{H}-E$ for the Hamiltonian. Figure 1 illustrates this structural equivalence: SVMs invert $\mathbf{Q}$ to obtain $\boldsymbol{\alpha}$, while quantum/PDE systems invert their Hamiltonian or differential operator to obtain Green’s function $G$.
As stated previously, a rigorous equivalence between kernels and Green’s functions demands that the following conditions are met:
@@
3.2.2 Eigenfunction vs. PDE expansion: Mercer’s theorem
Mercer’s theorem [13] originally emerged from the study of integral equations and provided a rigorous mathematical foundation for symmetric positive-definite kernels. The theorem’s relevance for machine learning was established through its application to kernel methods [11, 3]. This connection enabled the use of kernels in SVMs and beyond, thereby pushing forward the field of machine learning by allowing it to embed data into high-dimensional feature spaces without explicitly constructing these mappings.
As for Green’s functions, the expansion of partial differential equations in terms of eigenfunctions [14] is a cornerstone in classical analysis. Mercer’s theorem and PDE expansion both involve the decomposition of operators into eigenfunctions, with Mercer kernels providing a spectral decomposition similar to the eigenfunction solutions of PDEs.
Both the Mercer expansion for a positive semi-definite kernel and the eigenfunction expansion of a self-adjoint operator’s Green’s function decompose the respective objects into a sum of rank-one contributions. This similarity underscores the fact that in both cases the underlying eigenvalues and eigenfunctions govern the behavior of the system.
Spectral or eigenfunction decompositions.
According to Mercer’s theorem and standard linear algebra, a symmetric positive semi-definite kernel $K$ can always be expressed as a spectral decomposition:
@@
In quantum mechanics or PDEs, the resolvent operator $(\hat{H}-E)^{-1}$ associated with a Hamiltonian $\hat{H}$ can similarly be expressed as:
@@
Both $K$ and $G$ share a common structure: they are sums of rank-1 outer products weighted by spectral coefficients. The kernel $K$ corresponds to a weighted inner product in a high-dimensional feature space, while $G$ represents the propagator in a quantum or PDE system.
@@
In machine learning, positive semi-definite (PSD) kernel functions serve as the cornerstone of SVMs and kernel methods. By implicitly mapping input data into high-dimensional feature spaces, kernel functions enable nonlinear separations while satisfying Mercer’s theorem to ensure positive definiteness [3, 11].
Kernel functions can approximate the Green’s functions of PDEs. For example: Eq. (23) is analogous to the spectral decomposition of Green’s functions in PDEs, as in Eq. (24). While the mathematical structures are similar, their application backgrounds differ. Kernel functions are primarily used to capture data relationships in high-dimensional spaces, while Green’s functions are fundamental solutions to physical equations.
Positivity.
For $K(\mathbf{x},\mathbf{x}^{\prime})$ to be a valid SVM kernel on domain $\Omega$, it must be continuous, symmetric, and positive semi-definite. This ensures a convex SVM dual. For $G(\mathbf{x},\mathbf{x}^{\prime})$ to serve as an inverse of $(\hat{H}-E)$, one typically requires self-adjointness and positivity. Boundary conditions in PDE contexts (Dirichlet, Neumann, etc.) are responsible for the exact form of $G$.
@@
does not universally guarantee positive semi-definiteness. Although included as a standard option in many SVM implementations, the sigmoid kernel only meets Mercer’s condition under restrictive parameter regimes (e.g., specific ranges of $\gamma>0$ and $r<0$). Outside these regimes, the kernel matrix may lose its positive semi-definiteness, leading to non-convex optimization or instability in training.
This sensitivity to parameter choices underlines the reason why the sigmoid kernel is less used in practice than kernels like the RBF that satisfy Mercer’s theorem. This also complicates its use in drawing rigorous parallels between SVM kernels and Green’s functions, since the latter rely on strict positivity and well-defined spectral properties. As a result, while attempting to establish a precise equivalence between SVM kernels and Green’s functions, one must either operate within the safe parameter bounds for the sigmoid kernel or consider alternative kernel choices that inherently satisfy the necessary positivity conditions.
3.2.3 Margin vs. action extremization
Both SVMs and quantum mechanics involve optimizing a functional. For SVMs, the margin is optimized through an objective that depends on a kernel $K$. In quantum and Langrangian mechanics, the action $S$ is extremized, which often leads to Green’s functions through operator inversion.
In SVM, the decision hyperplane, which separates classes or fits a regression function, arises from the corresponding optimization problem, which is often solved via its dual formulation:
@@
We can draw a parallel between SVM margin maximization and action extremization in physics as follows. In SVMs, one seeks to maximize the margin by solving a variational problem in a high-dimensional feature space, which leads to a solution expressible as a weighted sum of kernel evaluations. In a similar manner, physical systems are often characterized by the principle of least action, where the trajectory or state of the system is obtained by extremizing a Lagrangian functional. In both cases, the solution is determined by finding a stationary point of an objective functional, and - in view of the representer theorem - the optimal solution can be represented as a finite linear combination of basis functions (kernel functions in SVMs, or Green’s functions in physical systems).
@@
For this reason, we may interpret a physical system’s trajectory or field configuration obtained via the Lagrangian as analogous to the SVM’s decision hyperplane. Determination of the optimal trajectory through a variational principle resembles selecting the optimal hyperplane in SVMs via margin maximization.
This perspective allows us to recast certain equations of motion (e.g., in regression or classification tasks) within an SVM framework, particularly when the operator or Green’s function governing the physical system is used as the SVM kernel.
3.2.4 Boundary conditions
Boundary conditions play an important role in defining Green’s functions as they uniquely determine the solution to the associated differential operator. Establishing a rigorous equivalence between Green’s functions and SVM kernels would, in principle, require a careful alignment of boundary conditions. However, this alignment is often difficult to prove in practice. Verifying that boundary conditions are met or, at the least, appropriately mirrored in the kernel construction is nontrivial and often infeasible for complex data-driven applications.
In this work, we therefore take a pragmatic approach. We don’t aim for a strict equivalence with respect to boundary conditions. Instead, we focus on the structural and mathematical form of the Green functions and their correspondence to kernel functions. Then, we demonstrate numerically that this relaxed equivalence is sufficient to yield effective kernel functions in practice. The empirical performance of Green-function-inspired kernels indicates that, even without rigorous boundary condition matching, these kernels can capture meaningful structure in data. This supports the idea that the mathematical form, rather than exact physical conditions, may be the more critical factor in many machine learning applications.
4 SVM regression in physical systems
In this section, we consider several aspects of physical systems and how particular kernels perform for these systems. All the examples involve regression, which is a standard application for SVMs [30].
The explored applications of SVM regression in modeling physical systems include electrical conductivity, scattering amplitudes, anharmonic oscillator energy levels, and photonic crystals.
For physical modeling of conductivity we will be using the Kubo formula. This procedure is well-documented: see, for instance, Mahan [31]. In electron-proton scattering, we apply the Born approximation, the description of which can be readily found in standard quantum mechanics references [33].